Find the optimum value of the function f) = 1208-45x + 4088 +5. state of the function attains a minimum on a maximum point
Answer #1

Given function - y = f(hi = 12 n5 - usn +4073 + 5 * = 60%- 18093 +12012 to dn for maximum / Minimum point fut az = 0 dn 60 nthan on n = o, o, 1, 2 (4 point obtationed To check maxima/or minima for particular point find 2 . dy = 240p3 - 54042 +24on aday = 240 (1) 3 - 540 (11² + 240 (1) ana = -60 dy co than Point n=1 is show maxima point dne = 240 (2) 3 - 540(21² +240x2 I +Now find value of function 191 neoo Elnas = 12 *(1) 5 - 450014 +40 (0)3 +5 = 12-45+40+5 = 12 21n-2 = 12 (2)5 - 45 (214 + 40 (